Two point charges fixed at some distance having ratio of magnitude k>1 and force between them is f1 if total charge is divided between them the orce becomes f2. If f1:f2=21:4 find k ?

Grade:12

1 Answers

Vikas TU
7 years ago
let q1 and q2 be the charges and the ratio is k.
Therfore,
q1/q2 = k 
or
q1 = k*q2
f1 = kq1q2/r^2
and
f2 = k((q1+q2)/2)^2/r^2
f1/f2 = 21/4 = 4*q1*q2/(q1 +q2)^2
substutute q1= kq2 we get,
k = 21*25/64 => 8.203